The base diameter and altitude of a right circular cone are observed at a certain instant to be 10 and 20 inches, respectively. If the lateral area is constant and the base diameter is increasing at a rate of 1 inch per minute, find the rate at which the altitude is decreasing.


1
Expert's answer
2021-01-21T15:04:40-0500

"Solution: Let ~r ~be~ diameter~ and~ h~ be~ the~ altitude~ of~ the~ right~ circular ~~cone.\n\\\\Then~ lateral ~area ~of~ right ~circular~ cone~ is ~=A=\u03c0r\\sqrt{h^2+r\u200b\u200b\u200b\u200b\u200b\u200b^r}"

"Differentiate ~~implicitly~ with ~respect~ to ~t\n\\\\\\frac{d}{dt}(A)=\\frac{d}{dt}(\u03c0r\\sqrt{h^2+r\u200b\u200b\u200b\u200b\u200b\u200b^r})"

"\Rightarrow \frac{dA}{dt}=\frac{\pi}{2}[\frac{r^2}{2}\frac{1}{2}\sqrt{(\frac{r^2}{2}+h^2)(r\frac{dr}{dt}+2h\frac{dh}{dt})}+\sqrt{\frac{r^2}{2}+h^2r\frac{dr}{dt}}] ...................................(I)"

"Given~that~\frac{dA}{dt}=0. Also ~r=10~inches~and~h~=20~inches~and~ \frac{dr}{dt}=0"

"Put~ all ~values ~in ~equation~ (I),then ~we~ get,
\\351.87+74.09\frac{dh}{dt}=0
\\74.09\frac{dh}{dt}=-351.87
\\\frac{dh}{dt}=-\frac{351.87}{74.09}
\\\frac{dh}{dt}=-4.75
\\Negative ~sign~ indicates~ that~ altitude ~is ~decreasing.
\\Thus,~the~ rate ~at~ which ~altitude ~is~ decreasing ~be ~4.75~ per ~minute."